# Calibration and Option Pricing with stochastic volatility and double exponential jumps

### v2

@@ 1.3. Key contributions
  • Pricing and analysis for several generations of exotic contracts (Asian, Discrete Variance Swap, Cliquet, Barrier).
@@ 2.1. Model specification
The real number $p$ (resp. $1-p$) represents the probability of jumping upward (resp. downward); conditionally on jumping upward (resp. downward), the size of jump follows an exponential distribution of parameter $\eta_{1}$ (resp. $\eta_{2}$). Note that the asymmetry induced by $\eta_{1},\eta_{2}$ and $p$ allows HKDE to capture more complex phenomena than Bates (where the jump size distribution is symmetric). Not only is this preferred from an empirical perspective (matching the asymmetry generally found in asset price jumps), but it provides additional flexibility to the model when calibrating to the observed market for derivatives. Last, the process $N:=(N_{t})_{t\geq 0}$ is a Poisson process with rate $\lambda>0$, and the volatility process $V:=(V_{t})_{t\geq 0}$ is a CIR process with mean $\theta$, speed of mean-reversion $\kappa$ and volatility of volatility $\sigma_{v}$.
@@ 2.2. Characteristic function
×$\times\exp(\eta\kappa\lambda^{-2}((\kappa-i\xi\lambda\rho-d(\xi))t-2\log((1-g(\xi)e^{-d(\xi)t})/(1-g(\xi)))))$
×$\times\exp(V_{0}^{2}\lambda^{-2}(\kappa-i\xi\lambda\rho-d(\xi))(1-e^{-d(\xi)t})/(1-g(\xi)e^{-d(\xi)t})),$
@@
$\begin{cases}d(\xi)\coloneqq\sqrt{(i\xi\lambda-\kappa)^{2}+\lambda^{2}(i\xi+\xi^{2})},\\ \displaystyle g(\xi)\coloneqq\frac{\kappa-i\xi\lambda\rho-d}{\kappa-i\xi\lambda\rho+d}.\end{cases}$
@@
(4) $\quad\phi_{\mathrm{Kou}}(\xi,t)=\exp\left(\lambda t\left(\frac{p\eta_{1}}{\eta_{1}-i\xi}+\frac{(1-p)\eta_{2}}{\eta_{2}+i\xi}-1\right)\right).$
@@ 2.3. Cumulants
$\kappa_{\mathrm{Kou},1}(t)$ $=t\lambda\left(\frac{p}{\eta_{1}}-\frac{1-p}{\eta_{2}}\right),$
$\kappa_{\mathrm{Kou},2}(t)$ $=2t\lambda\left(\frac{p}{\eta_{1}^{2}}+\frac{1-p}{\eta_{2}^{2}}\right),$
$\kappa_{\mathrm{Kou},3}(t)$ $=6t\lambda\left(\frac{p}{\eta_{1}^{3}}-\frac{1-p}{\eta_{2}^{3}}\right),$
$\kappa_{\mathrm{Kou},4}(t)$ $=24t\lambda\left(\frac{p}{\eta_{1}^{4}}+\frac{1-p}{\eta_{2}^{4}}\right).$
@@ 2.4. Parameters Analysis
In order to gain more intuition on the HKDE model, we show how the parameters influence the (spot) implied volatility surface. Specifically, we begin with a set of parameters $\Theta=(V_{0},\theta,\kappa,\sigma_{v},\rho,\lambda,p,\eta_{1},\eta_{2})$, and independently modify each parameter of $\Theta$ to observe the impact on the smile’s level, shape, convexity, etc.
@@ 2.4.1. Volatility parameters
(a) Bumping $\theta$ maturity 3 months.
(b) Bumping $\theta$ maturity 6 months.
(c) Bumping $\kappa$ maturity 3 months.
(d) Bumping $\kappa$ maturity 6 months.
@@
(a) Bumping $\kappa$ Effects on the term structure.
(b) Bumping $\theta$ Effects on the term structure.
@@ 2.4.2. Jump parameters
(a) Bumping $\sigma_{v}$ maturity 3 months.
(b) Bumping $\sigma_{v}$ maturity 6 months.
@@
(a) Bumping $\lambda$ maturity 3 months.
(b) Bumping $\lambda$ maturity 6 months.
@@
(a) Bumping $\eta_{1}$ maturity 3 months.
(b) Bumping $\eta_{1}$ maturity 6 months.
(c) Bumping $\eta_{2}$ maturity 3 months.
(d) Bumping $\eta_{2}$ maturity 6 months.
@@
First, we note (unsurprisingly) that increasing the jump intensity $\lambda$, which controls the frequency of observed jumps, shifts the smile upward for all levels of moneyness (see Figure 4). Second and more interestingly, the jump parameters $\eta_{1}$ and $\eta_{2}$ have a more significant and subtle impact on the shape of the implied volatility smile. Recall that the smaller the parameter $\eta_{1}$, the greater the magnitude of positive jumps; this increase in the jump size triggers a significant increase in the implied volatility for near ATM and ITM options (see Figure 5). The same effect holds for negative jumps driven by the parameter $\eta_{2}$. However, because the calibrated value of $p$ from Table 1 implies a much higher probability of positive jumps ($p=0.958$), the impact of negative jumps is more moderate compared to positive ones. It is therefore clear that the presence of two distinct tail heaviness parameters allows the HKDE model to capture more information from deep ITM and OTM prices, by adapting each wing of the implied volatility smile independently. This constitutes a clear improvement compared to the Bates model, where the symmetric distribution of jump size (log-normal) does not allow for a proper generation of asymmetric wings. As the wings are typically the hardest feature of the smile for a model to match, this places additional pressure on the other parameters of the Bates model to match the market smiles.
@@ 3.1. Pricing via PROJ method
where $(\widetilde{\varphi}_{a,n})_{n\in\mathbb{Z}}$ is the dual basis of $(\varphi_{a,n})_{n\in\mathbb{Z}}$. It is worth noting that $(\widetilde{\varphi}_{a,n})_{n\in\mathbb{Z}}$ and $(\varphi_{a,n})_{n\in\mathbb{Z}}$ are bi-orthogonal, i.e. $\langle\widetilde{\varphi}_{a,n},\varphi_{a,m}\rangle=\delta_{n,m}$, and that, by the asymptotic density of $(\varphi_{a,n})_{n\in\mathbb{Z}}$, $||\bm{\pi}_{a}^{\perp}f-f||\to_{a\to 0}0$. The projection coefficients $\langle f,\widetilde{\varphi}_{a,n}\rangle$ are available in closed form:
@@
since $\varphi_{a,n}$ is supported on $[x_{n-1},x_{n+1}]$.
@@
In numerical pricing, we set a leftmost boundary grid point $x_{1}$, and a gridwidth of $2\bar{\alpha}$ is chosen to encompass the support of $f$. A number $N$ of basis elements is then selected, which controls the approximation accuracy, and restricts the span of the basis elements to the finite set $\{\varphi_{a,n}\}_{n=1}^{N}$, covering $[x_{1},x_{1}+2\bar{\alpha}]$.
@@ 3.2. Calibration Methodology
(12) $\bm{\theta}^{*}\coloneqq\arg\min_{\theta\in\Theta}\left[\sum_{n=1}^{N}\sum_{j=1}^{N_{n}}w_{j}^{n}\left(\mathcal{V}_{j}^{(n)}(\theta)-v_{j}^{(n)}\right)^{2}\right],$
where $\mathcal{V}_{j}^{(n)}(\theta)$ denotes the model price of an option with maturity $\mathcal{T}_{n}$ and strike $K_{j}^{(n)}$, under the set of parameters $\theta$. We use the well-known vega-weighting (see for example [16]):
(13) $w_{j}^{n}:=\left(S_{0}\psi(d_{1})\sqrt{\mathcal{T}_{n}}\right)^{-1},\quad d_{1}:=\frac{1}{\sigma_{j}^{(n)}\sqrt{\mathcal{T}_{n}}}\left(\log\frac{S_{0}}{K_{j}^{(n)}}+\mathcal{T}_{n}\left(r_{n}-q_{n}+\frac{1}{2}\left(\sigma_{j}^{(n)}\right)^{2}\right)\right),$
@@ 5.1. Recursive Return Derivatives
(14) $Y_{1}:=w_{M}\cdot h(R_{M})+\varrho_{M}$
$Y_{m}:=w_{M-(m-1)}\cdot h(R_{M-(m-1)})+g(Y_{m-1})+\varrho_{M-(m-1)},\quad m=2,\ldots,M,$
@@ 5.1.1. Asian Option Prices
The Asian payoff (due to its averaging) is clearly smoother than the other claims mentioned in this article, resulting in option prices that are less sensitive to the underlying model. In particular, capturing the overall volatility level (especially near the money) is more important than the surface wings (i.e. the tails of the transition density), resulting in less variation between different models, even across different model families. Consequently, the variance of the option prices is tightly contained across all underlyings. The highest percentage differences are consistently observed for OTM options. The most significant discrepancy is noted for the stock SHOP between the Heston and HKDE models, amounting to \$$\$0.41$, which represents a $13\%$ difference. However, as a consequence we can expect tighter bid-ask spreads on these contracts, and even a $13\%$ difference in price could be significant, depending on the market.
@@
Model ITM (\$currency-dollar\\$\$) ATM (\$currency-dollar\\$\$) OTM (\$currency-dollar\\$\$) ITM (\$currency-dollar\\$\$) ATM (\$currency-dollar\\$\$) OTM (\$currency-dollar\\$\$)
@@
Model ITM (\$currency-dollar\\$\$) ATM (\$currency-dollar\\$\$) OTM (\$currency-dollar\\$\$) ITM (\$currency-dollar\\$\$) ATM (\$currency-dollar\\$\$) OTM (\$currency-dollar\\$\$)
@@ 5.1.2. Discrete Variance Options Prices
Discrete Variance swaps have been priced using the methodology from [22], and the results are presented in Table 5. We provide the prices for Discrete Variance options at three strike levels: $K\in\{0.01,0.03,0.05\}$. The prices are given to the third decimal place to highlight the differences between the models.
For all tickers, we observe the highest differences, in percentage, for the strike $K=0.05$. This derivative exhibits significant sensitivity to the underlying model. The underlying SPOT exhibits the most significant discrepancies, showing a \$$\$0.094$ difference between Bates and BGM, which is a significant $81\%$ difference. We can observe that of the four model prices for SPOT, the Bates model is the clear outlier. Following SPOT, SHOP shows notable variations, with a $37\%$ difference between HKDE and Heston. This example clearly highlights the importance of proper model calibration, as well as the need to use multiple models to assess the reasonableness of calibrated prices. We also note that a Lévy model such as BGM should not be seriously considered for pricing realized-variance derivatives, as it has no mechanism to model the “vol-of-vol”, and is thus ill-suited for capturing and risk-managing variance sensitive products. By contrast, stochastic volatility models such as Heston and its jump extensions were specifically designed for this purpose.
@@
Table 5. Prices (\$)currency-dollar(\\$)( \$ ) for Variance Call options, computed using calibrated parameters.
@@ 5.1.3. Cliquet Options Prices
The prices for Cliquet options are detailed in Table 6, priced using the methodology in [20]. We present the prices for Cliquet options at three strike levels: $K\in\{0.5,1,1.5\}$. Similarly to Discrete Variance swaps, prices are shown to the third decimal place to emphasize the distinctions between the models. As anticipated, price differences for Cliquet options are more moderate compared to Variance Swaps, as the cap and floors effectively limit the extent of tail differences between the models. This is of course dependent on how wide/symmetric the caps/floor are set. The most significant price variations are observed in SHOP and AMZN. For both tickers, the highest price differences occur at the strike $K=1.5$, with the Heston model providing the lowest prices. Specifically, in SHOP, the HKDE model shows the highest price difference, with a 17% spread between the models, while in AMZN, the BGM model shows a 15% spread. The observation that Heston produces lower prices than the models with a jump component highlights the importance of modeling jumps for return-sensitive assets such as cliquets, which can be very sensitive to the occurrence of price jumps. As the contracts “reset” with each monitoring point, it allows for multiple jumps during the life of the contract to produce outsized returns, as the cap permits.
@@
Table 6. Prices (\$)currency-dollar(\\$)( \$ ) for Cliquet options, computed using calibrated parameters.
@@ 5.2.1. Barrier Option Prices
Model ITM (\$currency-dollar\\$\$) ATM (\$currency-dollar\\$\$) OTM (\$currency-dollar\\$\$) ITM (\$currency-dollar\\$\$) ATM (\$currency-dollar\\$\$) OTM (\$currency-dollar\\$\$)
@@
Model ITM (\$currency-dollar\\$\$) ATM (\$currency-dollar\\$\$) OTM (\$currency-dollar\\$\$) ITM (\$currency-dollar\\$\$) ATM (\$currency-dollar\\$\$) OTM (\$currency-dollar\\$\$)
@@
Spot = \$100, ITM strike = \$70, OTM strike = \$130, upper barrier = \$$\$140$, maturity = 1 year ($T=1)$, monitoring points = 40, risk-free rate = $5\%$, dividend yield = $0\%$.

### v3

@@ 1.3. Key contributions
  • Pricing and analysis for several generations of exotic contracts (Asian, Discrete Variance Options, Cliquet, Barrier).
@@ 2.1. Model specification
The real number $p$ (resp. 1p1-p) represents the probability of jumping upward (resp. downward); conditionally on jumping upward (resp. downward), the size of jump follows an exponential distribution of parameter $\eta_{1}$ (resp. $\eta_{2}$). Note that the asymmetry induced by $\eta_{1},\eta_{2}$ and $p$ allows HKDE to capture more complex phenomena than Bates (where the jump size distribution is symmetric). Not only is this preferred from an empirical perspective (matching the asymmetry generally found in asset price jumps), but it provides additional flexibility to the model when calibrating to the observed market for derivatives. Last, the process $N:=(N_{t})_{t\geq 0}$ is a Poisson process with rate $\lambda>0$, and the volatility process $V:=(V_{t})_{t\geq 0}$ is a CIR process with mean $\theta$, speed of mean-reversion $\kappa$ and volatility of volatility $\sigma_{v}$.
@@ 2.2. Characteristic function
$\times\exp(\eta\kappa\sigma_{v}^{-2}((\kappa-i\xi\sigma_{v}\rho-d(\xi))t-2\log((1-g(\xi)e^{-d(\xi)t})/(1-g(\xi)))))$
$\times\exp(V_{0}^{2}\sigma_{v}^{-2}(\kappa-i\xi\sigma_{v}\rho-d(\xi))(1-e^{-d(\xi)t})/(1-g(\xi)e^{-d(\xi)t})),$
@@
$\begin{cases}d(\xi)\coloneqq\sqrt{(i\xi\sigma_{v}-\kappa)^{2}+\sigma_{v}^{2}(i\xi+\xi^{2})},\\ \displaystyle g(\xi)\coloneqq\frac{\kappa-i\xi\sigma_{v}\rho-d(\xi)}{\kappa-i\xi\sigma_{v}\rho+d(\xi)}.\end{cases}$
@@
(4) $\quad\phi_{\mathrm{Kou}}(\xi,t)=\exp\left(t\left(\lambda\left(\frac{p\eta_{1}}{\eta_{1}-i\xi}+\frac{(1-p)\eta_{2}}{\eta_{2}+i\xi}-1\right)+\omega_{\mathrm{KDE}}\cdot i\xi\right)\right),$
where the term $\omega_{\mathrm{KDE}}$ ensures that the combined model is risk-neutral, and is defined as follows:
$\omega_{\mathrm{KDE}}:=-\lambda\Big{(}\frac{p\eta_{1}}{\eta_{1}-1}+\frac{(1-p)\eta_{2}}{\eta_{2}+1}-1\Big{)}.$
@@ 2.3. Cumulants
$\kappa_{\mathrm{Kou},1}(t)$ $=t\left(\lambda\left(\frac{p}{\eta_{1}}-\frac{1-p}{\eta_{2}}\right)+\omega_{\mathrm{KDE}}\right),$
$\kappa_{\mathrm{Kou},n}(t)$ $=n!t\lambda\left(\frac{p}{\eta_{1}^{n}}+(-1)^{n}\frac{1-p}{\eta_{2}^{n}}\right),\quad\forall n\geq 2.$
@@ 2.4. Parameters Analysis
In order to gain more intuition on the HKDE model, we show how the parameters influence the (spot) implied volatility surface. Specifically, we begin with a set of parameters $P=(V_{0},\theta,\kappa,\sigma_{v},\rho,\lambda,p,\eta_{1},\eta_{2})$, and independently modify each parameter of $P$ to observe the impact on the smile’s level, shape, convexity, etc.
@@ 2.4.1. Volatility parameters
(a) Bumping $\theta$ - maturity 3 months.
(b) Bumping $\theta$ - maturity 6 months.
(c) Bumping $\kappa$ - maturity 3 months.
(d) Bumping $\kappa$ - maturity 6 months.
@@
(a) Bumping $\kappa$ - Effects on the term structure.
(b) Bumping $\theta$ - Effects on the term structure.
@@ 2.4.2. Jump parameters
(a) Bumping $\sigma_{v}$ - maturity 3 months.
(b) Bumping $\sigma_{v}$ - maturity 6 months.
@@
(a) Bumping $\lambda$ - maturity 3 months.
(b) Bumping $\lambda$ - maturity 6 months.
@@
(a) Bumping $\eta_{1}$ - maturity 3 months.
(b) Bumping $\eta_{1}$ - maturity 6 months.
(c) Bumping $\eta_{2}$ - maturity 3 months.
(d) Bumping $\eta_{2}$ - maturity 6 months.
@@
First, we note (unsurprisingly) that increasing the jump intensity $\lambda$, which controls the frequency of observed jumps, shifts the smile upward for all levels of moneyness (see Figure 4). Second and more interestingly, the jump parameters $\eta_{1}$ and $\eta_{2}$ have a more significant and subtle impact on the shape of the implied volatility smile. Recall that the smaller the parameter $\eta_{1}$, the greater the magnitude of positive jumps; this increase in the jump size triggers a significant increase in the implied volatility for near ATM and OTM call options (see Figure 5). The same effect holds for negative jumps driven by the parameter $\eta_{2}$. However, because the calibrated value of $p$ from Table 1 implies a much higher probability of positive jumps ($p=0.958$), the impact of negative jumps is more moderate compared to positive ones. It is therefore clear that the presence of two distinct tail heaviness parameters allows the HKDE model to capture more information from deep ITM and OTM prices, by adapting each wing of the implied volatility smile independently. This constitutes a clear improvement compared to the Bates model, where the symmetric distribution of jump size (log-normal) does not allow for a proper generation of asymmetric wings. As the wings are typically the hardest feature of the smile for a model to match, this places additional pressure on the other parameters of the Bates model to match the market smiles.
@@ 3.1. Pricing via PROJ method
where $(\widetilde{\varphi}_{a,n})_{n\in\mathbb{Z}}$ is the dual basis of $(\varphi_{a,n})_{n\in\mathbb{Z}}$. It is worth noting that $(\widetilde{\varphi}_{a,n})_{n\in\mathbb{Z}}$ and $(\varphi_{a,n})_{n\in\mathbb{Z}}$ are bi-orthogonal, i.e. $\langle\widetilde{\varphi}_{a,n},\varphi_{a,m}\rangle=\delta_{n,m}$, and that, by the asymptotic density of $(\varphi_{a,n})_{n\in\mathbb{Z}}$, $\bm{\pi}_{a}^{\perp}f-f||\to_{a\to 0}0$. The projection coefficients $\langle f,\widetilde{\varphi}_{a,n}\rangle$ are available in closed form:
@@
since $\varphi_{a,n}$ is supported on $x_{n-1},x_{n+1}]$.
@@
In numerical pricing, we set a leftmost boundary grid point $x_{1}$, and a gridwidth of $2\bar{\alpha}$ is chosen to encompass the support of $f$. A number $N$ of basis elements is then selected, which controls the approximation accuracy, and restricts the span of the basis elements to the finite set $\{\varphi_{a,n}\}_{n=1}^{N}$, covering $x_{1},x_{1}+2\bar{\alpha}]$.
@@ 3.2. Calibration Methodology
(12) $\bm{P}^{*}\coloneqq\arg\min_{P}\left[\sum_{n=1}^{N}\sum_{j=1}^{N_{n}}w_{j}^{(n)}\left(\mathcal{V}_{j}^{(n)}(P)-v_{j}^{(n)}\right)^{2}\right],$
where $\mathcal{V}_{j}^{(n)}(P)$ denotes the model price of an option with maturity $\mathcal{T}_{n}$ and strike $K_{j}^{(n)}$, under the set of parameters $P$. We use the well-known vega-weighting (see for example [16]):
(13) $w_{j}^{(n)}:=\left(S_{0}\psi(d_{1})\sqrt{\mathcal{T}_{n}}\right)^{-1},\quad d_{1}:=\frac{1}{\sigma_{j}^{(n)}\sqrt{\mathcal{T}_{n}}}\left(\log\frac{S_{0}}{K_{j}^{(n)}}+\mathcal{T}_{n}\left(r_{n}-q_{n}+\frac{1}{2}\left(\sigma_{j}^{(n)}\right)^{2}\right)\right),$
@@ 5.1. Recursive Return Derivatives
(14) $Y_{1}=w_{M}\cdot h(R_{M})+\varrho_{M}$
$Y_{m}=w_{M-(m-1)}\cdot h(R_{M-(m-1)})+g(Y_{m-1})+\varrho_{M-(m-1)},\quad m=2,\ldots,M,$
@@ 5.1.1. Asian Option Prices
The Asian payoff (due to its averaging) is clearly smoother than the other claims mentioned in this article, resulting in option prices that are less sensitive to the underlying model. In particular, capturing the overall volatility level (especially near the money) is more important than the surface wings (i.e. the tails of the transition density), resulting in less variation between different models, even across different model families. Consequently, the variance of the option prices is tightly contained across all underlyings. The highest percentage differences are consistently observed for OTM options. The most significant discrepancy is noted for the stock SHOP between the Heston and HKDE models, amounting to \$0.41\\$0.41, which represents a $13\%$ difference. However, as a consequence we can expect tighter bid-ask spreads on these contracts, and even a $13\%$ difference in price could be significant, depending on the market.
@@
Model ITM (\$\\$) ATM (\$\\$) OTM (\$\\$) ITM (\$\\$) ATM (\$\\$) OTM (\$\\$)
@@
Model ITM (\$\\$) ATM (\$\\$) OTM (\$\\$) ITM (\$\\$) ATM (\$\\$) OTM (\$\\$)
@@ 5.1.2. Discrete Variance Options Prices
Discrete Variance options have been priced using the methodology from [22], and the results are presented in Table 5. We provide the prices for Discrete Variance options at three strike levels: $K\in\{0.01,0.03,0.05\}$. The prices are given to the third decimal place to highlight the differences between the models.
For all tickers, we observe the highest differences, in percentage, for the strike $K=0.05$. This derivative exhibits significant sensitivity to the underlying model. The underlying SPOT exhibits the most significant discrepancies, showing a \$0.094\\$0.094 difference between Bates and BGM, which is a significant $81\%$ difference. We can observe that of the four model prices for SPOT, the Bates model is the clear outlier. Following SPOT, SHOP shows notable variations, with a $37\%$ difference between HKDE and Heston. This example clearly highlights the importance of proper model calibration, as well as the need to use multiple models to assess the reasonableness of calibrated prices. We also note that a Lévy model such as BGM should not be seriously considered for pricing realized-variance derivatives, as it has no mechanism to model the “vol-of-vol”, and is thus ill-suited for capturing and risk-managing variance sensitive products. By contrast, stochastic volatility models such as Heston and its jump extensions were specifically designed for this purpose.
@@
Table 5. Prices (\$)(\\$) for Variance Call options, computed using calibrated parameters.
@@ 5.1.3. Cliquet Options Prices
The prices for Cliquet options are detailed in Table 6, priced using the methodology in [20]. We present the prices for Cliquet options at three strike levels: $K\in\{0.5,1,1.5\}$. Similarly to Discrete Variance options, prices are shown to the third decimal place to emphasize the distinctions between the models. As anticipated, price differences for Cliquet options are more moderate compared to Variance Options, as the cap and floors effectively limit the extent of tail differences between the models. This is of course dependent on how wide/symmetric the caps/floor are set. The most significant price variations are observed in SHOP and AMZN. For both tickers, the highest price differences occur at the strike $K=1.5$, with the Heston model providing the lowest prices. Specifically, in SHOP, the HKDE model shows the highest price difference, with a 17% spread between the models, while in AMZN, the BGM model shows a 15% spread. The observation that Heston produces lower prices than the models with a jump component highlights the importance of modeling jumps for return-sensitive assets such as cliquets, which can be very sensitive to the occurrence of price jumps. As the contracts “reset” with each monitoring point, it allows for multiple jumps during the life of the contract to produce outsized returns, as the cap permits.
@@
Table 6. Prices (\$)(\\$) for Cliquet options, computed using calibrated parameters.
@@ 5.2.1. Barrier Option Prices
Model ITM (\$\\$) ATM (\$\\$) OTM (\$\\$) ITM (\$\\$) ATM (\$\\$) OTM (\$\\$)
@@
Model ITM (\$\\$) ATM (\$\\$) OTM (\$\\$) ITM (\$\\$) ATM (\$\\$) OTM (\$\\$)
@@
Spot = \$100, ITM strike = \$70, OTM strike = \$130, upper barrier = \$140\\$140, maturity = 1 year ($T=1)$, monitoring points = 40, risk-free rate = $5\%$, dividend yield = $0\%$.